# The Complexity Dynamics of Grokking

### v1

@@ 5 Experiments
To study the complexity dynamics of networks transitioning from memorizing to generalizing solutions, we adopt the “grokking" tasks first reported in Power et al. (2022). These are a set of modular arithmetic tasks which networks easily over-fit, quickly achieving zero training error while test error remains large, suggesting memorization of the training examples. After many further steps of gradient descent after over-fitting, networks suddenly generalize and achieve perfect test accuracy. Although grokking is a sign of improperly regularized training, the extended delay between memorization and generalization is convenient for understanding complexity dynamics, and their relationship with generalization.
@@
We reproduce a subset of the grokking experiments first reported in Power et al. (2022). Our training tasks all consist of learning a binary operation mod a prime number $p=113$. For each binary operation we construct a dataset of equations of the form $\langle x\rangle\langle\text{op}\rangle\langle y\rangle\langle=\rangle\langle x\circ y\rangle$, where $\langle a\rangle$ stands for the token corresponding to element $a$. We train standard decoder-only transformers with a single layer for mod addition and multiplication, and two layers for mod subtraction and division. Further training details can be found in the appendix. We use Algorithm 1 to compute complexity upper-bounds on the coarse-grained models. The final compression step is performed by bzip2, an off-the-shelf compression utility.
Figure 10 in the appendix is a sanity check to ensure that grokking occurs: indeed, the regularized models generalize, while the unregularized models never achieve generalization. Fig 2 plots accuracy and complexity together for our regularized models, and indicates that as we expected from Occam’s Razor, complexity is maximized during memorization, and falls during generalization. Next, we plot algorithmic rate–distortion curves of best fit in Fig 3. We fit curves of the form $y=a\log(x)+b$, and display $R^{2}$ values in the legend. These curves show how lossy model compression trades off with lossless data compression for the converged models at varying distortion bounds.
We compare the total description length of all models in Fig 4, and find that the models produced by our regularizer achieve the best explanation of the dataset amongst the methods. That is, they achieve a smaller sum of model complexity and data entropy vs the other methods, hence achieve the greatest compression of the data in this model class. Plotting the generalization bound from Eqn. 4, we see in Fig 5 that again, the models produced by our regularizer come closest to a non-trivial generalization bound. The goal of this work is to explain grokking, clarify the nature of complexity and its relationship to generalization, and provide a practical method of complexity approximation based on insights from lossy compression. In particular, we explained how lossy compression provides a principled approach to noise reduction and coarse-graining in networks. In particular, we did not attempt to achieve the tightest possible complexity bounds. Nevertheless, it is surprising that we get as close as we do to non-trivial generalization bounds, since the models we use are large, and the datasets small: $|D|=12769$ vs. $|D|\approx 1.5\times 10^{13}$ for contemporary LLMs (Dubey et al., 2024).
Figure 2: Complexity and accuracy vs steps for models regularized using our method. As models memorize, complexity increases. Complexity falls as generalization occurs. We use Algorithm 1 to coarse-grain models with $\epsilon=1$. Plots show the mean of six seeds with std. error shaded.
Figure 3: Algorithmic rate–distortion curves $r_{\theta}(\hat{\theta})$ for all datasets at varying distortion bounds $\epsilon$. The rate is the complexity of the coarse-grained weights $\hat{\theta}$ at the end of training. Mean over six seeds.
@@ 6 Conclusion
This work examines the complexity dynamics of neural networks, and proposes an explanation of the grokking phenomenon—the sudden transition from memorization to generalization long after the training data is over-fit. We do this by introducing a novel measure of intrinsic complexity based on Kolmogorov complexity and rate-distortion theory. We observe a consistent pattern of the rise and fall of complexity during training, corresponding to memorization followed by generalization. Our principled approach to lossy compression of neural networks formalizes the concept of "apparent complexity" and connects it to explicit generalization bounds. Our proposed spectral entropy regularization method encourages low-rank representations by penalizing a differentiable measure of the model’s effective dimension, which helps limit model complexity and enables tight compression bounds. We hope that a deeper view of complexity beyond mere parameter count and capacity measures will lead to greater insight into the nature of generalization in learning systems.
@@ Appendix A Appendix
[…]
@@
Modulus $p$ $113$
Dataset size $N$ $p^{2}=12769$
[…]
Transformer layers $n_{\text{layers}}$ (+,×:1),(,÷:2)(+$\times$:1),(-$\div$:2)( + , × : 1 ) , ( - , ÷ : 2 )
@@
Transformer head count $n_{\text{head}}$ $4$
@@
Noise regularization scale $\delta$ (+,×$(+,\times:10^{-2}),(-,\div:10^{-3})$
@@
Weight decay scalar - $1$

### v2

@@ 5 Experiments
To study the complexity dynamics of networks transitioning from memorizing to generalizing solutions, we adopt the “grokking” tasks first reported in [1]. These are a set of modular arithmetic tasks which networks easily over-fit, quickly achieving zero training error while test error remains large, suggesting memorization of the training examples. After many further steps of gradient descent after over-fitting, networks suddenly generalize and achieve perfect test accuracy. Although grokking is a sign of improperly regularized training, the extended delay between memorization and generalization is convenient for understanding complexity dynamics, and their relationship with generalization.
@@
We reproduce a subset of the grokking experiments first reported in [1]. Our training tasks all consist of learning a binary operation mod a prime number $p=113$. For each binary operation we construct a dataset of equations of the form $\langle x\rangle\langle\text{op}\rangle\langle y\rangle\langle=\rangle\langle x\circ y\rangle$, where $\langle a\rangle$ stands for the token corresponding to element $a$. We train standard decoder-only transformers with a single layer for mod addition and multiplication, and two layers for mod subtraction and division. Further training details can be found in the appendix. We use Algorithm 1 to compute complexity upper-bounds on the coarse-grained models. The final compression step is performed by bzip2, an off-the-shelf compression utility.
First, we establish that grokking occurs in all regularized models, and does not occur in unregularized models in Appendix Figure 10. Next, we plot accuracy and complexity together for the regularized models in Fig 2, which shows our main result: as expected from Occam’s Razor, complexity is maximized during memorization and falls during generalization, demonstrating a clear complexity phase transition. In contrast, Fig 3 demonstrates that without regularization, models tend to stay in the high-complexity memorization regime indefinitely. Next, we plot algorithmic rate–distortion curves of best fit in Fig 4. We fit curves of the form $y=a\log(x)+b$, and display $R^{2}$ values in the legend. These curves show how lossy model compression trades off with lossless data compression for the converged models at varying distortion bounds. Fig 5 shows how the complexity dynamics change under varying distortion levels $\epsilon$: we see that as the distortion bound is loosened (larger $\epsilon$), the corresponding complexity falls for the same model, as expected from the theory of rate–distortion.
Plotting the generalization bound from Eqn. 4, we see in Fig 6 the models trained with our regularizer come closest to a non-trivial generalization bound: they achieve the least sum of data entropy and model complexity. Fig 11 in the appendix plots the total description length of the dataset under each model, which mirrors the generalization bound.
@@ 6 Conclusion
Just as physical systems can transition between distinct phases of matter, we show that learning systems can transition between distinct phases characterized by the complexity of their internal representations. We suggest that a complexity phase transition is the central mechanism driving the grokking phenomenon, in contrast to prior works which focus on more specific signatures such as weight norms [10] or linear structure [8]. We justify our analysis on theoretical grounds by linking complexity to generalization through formal statistical generalization bounds [15]. From this perspective, we can see that small weight norms or linear structure are different ways for networks to be simple, but these metrics alone are insufficient explanations since complexity can emerge through a variety of mechanisms. For example, we showed how the effective rank of the network is one means by which complexity can emerge, which we penalize through the spectral entropy measure we introduced in Section 4.2.
The distinctive separation between the memorizing and generalizing phases makes the grokking phenomenon an ideal model system to initiate the study of complexity phase transitions in learning systems. A related pathological phenomenon observed in learning systems is double descent [12], in which test error first falls, then rises, then falls again as a function of model size, with training steps fixed. A unified theory of complexity and generalization should predict when and under which conditions such training dynamics occur, and may suggest avenues for further progress in the theory of machine learning. Generalizing our results to account for arbitrary learning systems could lead to a deeper understanding of the interplay between optimization dynamics, the loss landscape, data, model capacity, and complexity.
Equipped with such a theory, practitioners could predict the performance of machine learning systems before running costly experiments, and insights gained from a more complete theory of machine learning could enable new breakthroughs in our understanding of complex systems more broadly. The emerging field of algorithmic thermodynamics [29, 30, 31] provides a promising mathematical and computational framework to analyze these systems.
The study of phase transitions in learning systems provides a window into fundamental questions about the nature of generalization and the emergence of structure in dynamical systems. The rise and fall of complexity we report in this work is conjectured to be a generic property of certain complex dynamical systems [2], and understanding its origin in greater detail could help explain not only the emergence of abstraction and generalization in learning systems, but also the apparent complexity we observe in the natural world more broadly. Machine learning systems such as neural networks are convenient model complex systems to study from a theoretical perspective due to their transparency and controllability, and the study of their complexity-theoretic properties is a promising avenue for further research.
Acknowledgments
The authors thank Jasmine Brewer and Andrew P. Turner for valuable discussions that strengthened this work.
Funding
This work was supported by the EPSRC Programme Grant “From Sensing to Collaboration” (EP/V000748/1). Branton DeMoss was supported by an AWS Lighthouse Scholarship.
@@ Appendix A Appendix
 if $\left|\textsc{Loss}(\text{quantized\_params})-\textsc{Loss}(\text{params})\right|<\epsilon$ then
@@
Modulus pp $113$
Dataset size NN $p^{2}=12769$
Modular arithmetic tasks ${+,×,,÷}\{$$\times$$\div\} -$
Train data fraction per task - $(+,\times:20\%),(-,\div:30\%)$
Transformer layers $n_{\text{layers}}$ $(+,\times:1),(-,\div:2)$
@@
Transformer head count $n_{\text{head}}$ 44
@@
Noise regularization scale $\delta$ $(+,\times:10^{-2}),(-,\div:10^{-3})$
@@
Weight decay scalar - 11